find the dii\fferential equation of first order and first degree
dy/dx = sin (x+y)

Grade:12th pass

1 Answers

Aditya Gupta
4 years ago
the ques seems a bit wrongly worded. so i will answer what i understood it as. the order of diff eqn is 1 and degree is also 1.
now to solve we can put z=x+y so that dz/dx= 1+dy/dx= 1+sin (x+y)= 1+sinz
or dz/(1+sinz)= dx
integrate both sides 
C – 2/(1+tan(z/2)) = x
or 2/[1+tan((x+y)/2)] + x= C